Alb-Air is a low-cost airline that employs an average of 1,000 flight attendants on its aircraft. However, exhaustion and the very ordinary salary lead to a high turnover rate: once trained and on the job, flight attendants only stay with the company for an average of two years. Alb-Air must therefore train new employees regularly and economically. Following the mandatory six-week training, a new employee will take a week of paid vacation and will then be enrolled in the pool of available employees who will begin work as flight attendants leave after their average length of service. two years. Alb-Air pays new employees $ 125 per week while they are interns (6 week training + one-week vacation + time spent in the pool of available employees before taking up their position). A training session requires a team of ten instructors and ten assistants. The salary of an instructor is $ 220 / week and $ 80 / week for an assistant. The training team is only paid for the training periods it gives. In addition, the salary of instructors and assistants is independent of the number of new employees who attend a training session. It is also independent of the number of sessions offered at the same time. The other costs of hosting a training session are negligible. 1) Use the EOQ model to calculate the optimal size of a training session, the number of sessions to offer per year, and the time between each training session. Also estimate the total annual cost of this strategy. Note that in this context, this annual cost has three components, similar to the total inventory cost (orders + storage + purchase). Assume that there are 50 weeks in a year and that the number of applicants always exceeds the demand for flight attendants. 2) Modify your solution so that there is only one training session at a time. This means that a new session must start every six weeks. Calculate the size of these sessions and the total annual cost. Compare your result with that of the previous question.
Answer #1

1)As per the question mentioned, here we can use the method of EOQ(Economic Ordering Quantity) for solving this question.

EOQ = square root of(2 x A xO) / C

here the assumptions,

2 is constant figure

A is annual demand

O is ordering cost per order (it will be same whatever the lot of units ordered)

C carrying cost per unit.

in our solution we can substitute our properties in as,

A = 1000 attendants, O = $300 per week ($220+$80)*, C = $125 per week.**

subtituting this into the equation will be,

squar root of ( 2 x 1000 x 300) / 125 =69.28

which means 69.28 or 70 attendants will be optimum size.

as of this findings, the number of sessions comapny wants to offer to complete 1000 attendants will be

= 1000/70=14.28 sessions or 14.29 sessions.

if the company arranging 6 week training session as usual company can arrange 8.333 (50 weeks / 6 weeks) sessions per year. that means company will get 8.333 times of 6 week training session. but here company can accomodate mutitple sessions at the same time and applying EOQ shows that the 14.29 sessions of 70 attendants per session will be optimal. so ,

the number of sessions comapany wants to arrange in a 6 week training session =14.29/8.333=1.714 sessions per 6 weeks training time.which means one training can only take 3.5 weeks (6 weeks / 1.714 ).

annual cost will be

interns salary/week= $125 x 1000 = $125000

so per year of 50 weeks will be = 125000 x 50 =$6250000

salary of instructor and assistant per yaer=$300 x 50weeks = $15000

total =$6250000 + $15000 = $6265000

*&** if you are taking these figures as for one week or for 6 weeks of session the result will be same.

2)if the company cannot accomodate multiple training session at the same time,company can arrange as 120 (1000/50 x 6) attandatnts in a 6 week of session.

annual cost will remain same as the above solution. becasue whatever is company wants to train all the attendants in the year so the salary of the inerns will not change as a whole. and company wants to maintain the istructor for the 50 weeks ,whatever the company following pattern. so in the above both situation only the size of session will vary but the annual cost will not